# Understanding the differences in Foundation Models: Attention, State Space Models, and Recurrent Neural Networks

### v2

@@ 1 Introduction
Foundation models serve as the backbone for a wide range of tasks across Artificial Intelligence due to their ability to learn complex interactions in large datasets [Bommasani et al., 2021]. In recent years, the attention mechanism [Vaswani et al., 2017] has been the dominating token-mixing strategy in foundation models. However, its major computational bottleneck, i.e., the quadratic complexity with context length, has posed a challenge to scaling and deploying these models beyond moderate context lengths [Tay et al., 2021]. In order to mitigate these issues, attention-free architectures have been proposed: prominent examples of these are the novel State Space Models (SSMs) [Gu et al., 2022b, Smith et al., 2023, Orvieto et al., 2023, Gu and Dao, 2023, Dao and Gu, 2024], as well as recent efforts to enhance Recurrent Neural Networks (RNNs) [Stanić et al., 2023, De et al., 2024, Qin et al., 2024, Beck et al., 2024]. Although these models show great promise in boosting efficiency, current comparisons with attention are merely empirical. Despite the prevalence and ubiquity of foundation models, a principled understanding of the similarities and differences among these different design strategies is currently lacking.
In order to close this gap, we introduce the Dynamical Systems Framework (DSF), a theoretical framework that allows to evaluate the similarities and differences between different foundation models in a principled manner. This framework spans most current architectures and allows for direct comparisons, theoretical and computational, across attention, SSMs, and RNNs. The DSF provides new insights on the most relevant features found in current architectures, and can inform a systematic development of future hybrid models. Specifically, in this paper we answer the following questions:
  • How are attention, SSMs, and RNNs related?
    TL;DR: All three model classes can be represented as recurrent models that can directly be compared using the proposed DSF.
@@
  • How closely are linear attention and S6 (i.e. Mamba) related?
    TL;DR: The common feature is the coupling of state transition and input matrix via a single (normalization) parameter in recurrent representation. However, the two models differ in the parameterization of this parameter, which we analyze experimentally.
@@
Furthermore, it is important to highlight that, for the models studied here, some model classes are natively stated in recurrent form (i.e. SSMs, RNNs), while others are stated in convolutional (matrix) form (i.e. attention). The DSF allows to switch between these model classes and leverage computational tools developed for other classes. For instance, the recurrent form is efficiently implemented via scan algorithms [Blelloch, 1990], e.g., selective scan [Gu and Dao, 2023], parallel scan [Smith et al., 2023, Orvieto et al., 2023], and accelerated scan [Kyrylov, 2024]. The same holds for the convolutional form via, e.g., flash attention [Dao, 2023], flash linear attention [Yang and Zhang, 2024], and structured masked attention [Dao and Gu, 2024]. Given that the structural requirements on the model parameterization of the algorithm is met, the DSF allows to identify existing algorithms to apply to a new model even if the algorithm was designed for another model class.
@@
We use Latin letters in the following way: $N$ is the size of the hidden state in the DSF, $n$ the state expansion, $d$ the embedding size or model size, and $L$ the sequence length. A visual representation of these dimensions is given in Appendix A. We use superscripts, e.g. $\cdot^{d}$, to denote the elements or block-elements of a matrix and a block-matrix. We use subscripts, e.g. $\cdot_{i}$, to denote the time index (or input dependency). Specifically, $v_{i}$ represents the value of vector $v$ at time $i$. We use bold notation to indicate sequences, i.e., $\mathbf{v}_{i}=[v_{1},\dots,v_{i}]$. We use $\sigma(\cdot)$ to denote is the sigmoid function. The products $\odot$ and $\otimes$ denote the Hadamard (element-wise) product and the Kronecker (block-wise) product, respectively. $\mathbb{I}_{n}$ denotes the identity matrix of size $\mathbb{R}^{n\times n}$. Generally, we omit stating the bias term for weight matrices unless stating the bias term helps with clarity.
@@ 2 Preliminaries
In this section, we introduce the key architectural components studied in this work: attention, SSMs, and RNNs. We remark that these components are often the central block - considered to be the backbone - within a complex architecture composed of other blocks and skip connections (see for instance [Touvron et al., 2023]). In what follows, we review exclusively the backbone block, which we denote as $f(\cdot)$ in $\mathbf{y}=f(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{L\times d}$ and $\mathbf{y}\in\mathbb{R}^{L\times d}$ are the input and output sequences, respectively.
@@ 2.1 Attention
The standard self-attention block [Vaswani et al., 2017] consists of three matrices: $W_{Q},\ W_{K},\ \text{and}\ W_{V}$, which are the learnt parameters of the model. These matrices, when multiplied with the input $\mathbf{u}$, yield the queries $\mathbf{q}\in\mathbb{R}^{d_{k}}$, keys $\mathbf{k}\in\mathbb{R}^{d_{k}}$, and values $\mathbf{v}\in\mathbb{R}^{d_{v}}$, respectively:
@@
where $\zeta(\cdot)$ is a map $\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}$ and is applied row-wise. In the standard version of attention - softmax attention - $\zeta(\cdot):=\mathrm{softmax}(\cdot)$. However, given the limitations of the softmax function, alternative formulations have been proposed. In this work, we consider two formulations of attention: softmax attention (2) and linear attention [Katharopoulos et al., 2020]. We focus on the masked attention formulations, i.e., the attention matrix $\zeta(\mathbf{q}\mathbf{k}^{\top})$ has a lower-triangular structure. Furthermore, in order to simplify the derivations, we drop the scaling factor $\sqrt{d_{k}}$.
@@ 2.2 State Space Models
Architectures based on a state space parametrization compute the output $\mathbf{y}$ through a dynamic recurrence of input signals at each time step $i$,
@@
where $h_{i}$ is the hidden state of the system, and the dynamic matrices of appropriate dimensions $A_{i},B_{i},C_{i},D_{i}$ are the learnt parameters of the model. Different time-varying and time-invariant parameterizations for $A_{i},B_{i},C_{i},D_{i}$ have been proposed in the literature (an overview is given in [Amo Alonso et al., 2024]). Here we discuss the most prominent one.
@@
The first selective SSM parametrization (S6) was introduced together with the Mamba architecture [Gu and Dao, 2023]. The S6 block parametrizes the recurrence as
@@
with $\Delta_{i}=\mathrm{softplus}(W_{\Delta}(W_{u}u_{i})+b_{\Delta})$ for every $i$, $W_{\Delta}$, $W_{u}$, $W_{B}$, $W_{C}$, $W_{D},A$ are learnt matrices of appropriate dimensions, and $b_{\Delta}$ is a learnt bias. While SSM models allow for complex-valued matrices $A_{i},B_{i},C_{i},D_{i}$, here we restrict ourselves to real-valued matrices as in [Gu and Dao, 2023].
@@ 2.3 Recurrent Neural Networks
Similar to SSMs, RNNs also parameterize the input-output relationship via a recurrent computation, commonly given by the long short-term memory (LSTM) [Hochreiter and Schmidhuber, 1997], at each time step $i$
@@
where $W_{f},W_{i},W_{o}$ and $U_{f},U_{i},U_{o}$ are the learnt gate parameters. In this paper, we focus on two variants: quasi LSTMs (qLSTM) [Stanić et al., 2023], which removes the output dependence of the gates, and RG-LRU [De et al., 2024], which attempts to integrate ideas from SSMs into RNNs.
@@
The RG-LRU model presents a hybrid between a qLSTM and a SSM. The recurrence is given as
@@ 3.1 Dynamical Systems Framework (DSF)
The DSF relies on a dynamical systems representation of the architectures. As is standard in RNN and SSM literature, we choose a recurrent state space representation. In particular, a linear structured time-varying (LTV) dynamical system is defined by the recurrence
@@
where $h_{i}\in\mathbb{R}^{N}$ is the hidden state initialized with $h_{-1}=0$, $\Lambda_{i}\in\mathbb{R}^{N\times N}$ is the diagonal state transition matrix, $B_{i}\in\mathbb{R}^{N\times d}$ and $C_{i}\in\mathbb{R}^{d\times N}$ are the input and output matrices, respectively, and $D_{i}\in\mathbb{R}^{d\times d}$ is a scaled skip connection.
Dynamical system (11) can alternatively be written in its convolutional representation, i.e., $\mathbf{y}=\mathbf{\Phi}\mathbf{u}$, where the convolutional kernel $\mathbf{\Phi}$ is defined as
@@
Note that the convolution kernel $\mathbf{\Phi}\in\mathbb{R}^{L\times L}$ is of the same dimension as the attention matrix $\zeta(\mathbf{q}\mathbf{k}^{\top})$ and that these matrices are equivalent, up to the scaling factor $W_{V}$ used in self-attention.
@@
This recurrent view yields a causal convolution kernel by definition. However, certain models (e.g. non-masked attention) also use non-causal kernels. This can be incorporated in the DSF (11) by modifying the state update (11a). For the sake of simplicity and consistency with recent literature (RNNs, SSMs), we stick with causal models in the following.
@@ 3.2 Architecture Reformulation
In the following, we show how popular architectures based on attention, SSMs, and RNNs can be rewritten into the DSF. To do this, all models will be reformulated into recurrence (11), i.e., all resulting DSF representations will have hidden state dimension $N$ [^1] The dimensions used in the following are visualized in Appendix A for clarity. Although the parametrization of models commonly found in the literature is conductive to efficient computation, here we depart from this convention. The goal of the DSF reformulation is to establish a theoretical framework that leads us to mathematical insights on the design of these models. The presented formulations are not intended to be computationally implemented in DSF form, however the framework can be used to identify computational algorithms for new architectures as mentioned in the introduction.
@@ 3.2.1 Attention
In the following, we assume that we can separate the map in (2) as
@@
where $\phi(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, $\psi(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, and $\eta(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m\times(i+1)}\rightarrow\mathbb{R}$, which is the case for all the considered architectures in this paper. This allows us to write the self-attention input-output relationship as
@@
We note that for the recurrence (11), the matrix $\Lambda_{i}$ is given as an $nd\times nd$ matrix, where $n$ is the number of features in $\phi$ and $\psi$, and $d$ is the input dimension. However, due to the structure of $\Lambda_{i}$ as given in (15a), it can be implemented as a scalar multiplication $\frac{\eta(q_{i-1},\mathbf{k}_{i-1})}{\eta(q_{i},\mathbf{k}_{i})}h_{i-1}$ in (11). Hence, the hidden state is never materialized as such in the computation of the attention scores. Interested readers are referred to Appendix B for a detailed derivation.
@@
In the case of linear attention, both maps $\phi(\cdot)$ and $\psi(\cdot)$ in the DSF parametrization (15) are linear in their arguments and we use the kernel proposed in [Katharopoulos et al., 2020], i.e.,
@@
where elu stands for exponential linear unit.
@@
We also study here the generalized linear attention, where we simply require that $\phi(\cdot)$ and $\psi(\cdot)$ be linear in their arguments; but we do not require that the normalization function $\eta(q_{i},\mathbf{k}_{i})$ is linear as well.
@@
Softmax attention also satisfies the assumption of separability (13). However, it holds that the feature vector representation of the transformed Gaussian kernel in the softmax function, i.e., $e^{q_{i}^{\top}k_{j}}$, is infinite dimensional. Hence, the DSF representation (15) of softmax attention (2) and its corresponding hidden state dimension $N$ would also be infinite dimensional. This insight gives further motivation to approximations of the softmax function by using, e.g., a Taylor series approximation such as in [Nauen et al., 2024], to render the feature vector finite-dimensional.
@@
The exponential in the softmax attention $e^{q_{i}^{\top}k_{j}}$ can be expressed in terms of its Taylor expansion, which consists of an infinite sum of polynomial kernels of increasing degree $p$, decomposable through the vectors of monomials $\bigotimes_{j=1}^{p}q_{i}$. See Appendix C for a complete proof. ∎

### v3

@@ 1 Introduction
Foundation models serve as the backbone for a wide range of tasks across Artificial Intelligence due to their ability to learn complex interactions in large datasets [1]. In recent years, the attention mechanism [2] has been the dominating token-mixing strategy in foundation models. However, its major computational bottleneck, i.e., the quadratic complexity with context length, has posed a challenge to scaling and deploying these models beyond moderate context lengths [3]. In order to mitigate these issues, attention-free architectures have been proposed: prominent examples of these are the novel State Space Models (SSMs) [4, 5, 6, 7, 8], as well as recent efforts to enhance Recurrent Neural Networks (RNNs) [9, 10, 11, 12]. Although these models show great promise in boosting efficiency, efforts to provide a rigorous theoretical comparison are scarce, and current comparisons with attention are merely empirical (see Section 5 for an in-depth discussion). Despite the prevalence and ubiquity of foundation models, a principled understanding of the similarities and differences among these different design strategies is currently lacking.
In order to close this gap, we introduce the Dynamical Systems Framework (DSF) – a theoretical framework based on a control theoretic perspective – that allows us to evaluate the similarities and differences between different foundation models in a principled manner. The DSF serves as a powerful tool for approaching theoretical research questions about foundation models, enabling direct comparisons – both theoretical and experimental – across architectures such as attention mechanisms, SSMs, and RNNs. We believe that the DSF provides new insights on the most relevant features found in current architectures, and can inform a systematic development of future hybrid models. The DSF further simplifies identification of existing computational algorithms to apply to newly developed models. Rather than providing an exhaustive list of insights, the results included below are meant to exemplify important questions that the DSF can answer and guide future research. Specifically, we explore the following questions:
  • How are attention mechanisms, SSMs, and RNNs related?
    TL;DR: All three model classes can be represented as recurrent models, which can be compared using the proposed DSF.
@@
  • Why do SSMs significantly outperform attention on the LRA benchmark?
    TL;DR: The performance gap can be explained by the recurrent normalization strategy (discretization step) used by selective SSMs as discussed in Section 4.2.
  • How closely are linear attention, S6 (i.e. Mamba) related?
    TL;DR: The common feature is the coupling of state transition and input matrix via a single (normalization) parameter in their recurrent representation. However, the models differ in the parameterization of this parameter, which we analyze experimentally.
@@
We note that the main contribution of our paper is the introduction of the DSF, which is a unifying framework for analysis of attention mechanisms, SSMs, and RNNs. To the best of our knowledge, this is the first unified framework that allows analysis of all three model classes in the same parameterization and thus allows to identify differences in the models that lead to significant performance improvements. While some of the provided results already exist in the literature (e.g that increased state size improves performance), we also provide novel insights unique to the DSF framework in a comprehensive way that enables further analysis with control theoretical tools.
@@
We use Latin letters in the following way: $N$ is the size of the hidden state in the DSF, $n$ the state expansion, $d$ the embedding size or model size, and $L$ the sequence length. A visual representation of these dimensions is given in Appendix A. We use superscripts, e.g. $\cdot^{d}$, to denote the elements or block-elements of a matrix and a block-matrix. We use subscripts, e.g. $\cdot_{i}$, to denote the time index (or input dependency). Specifically, $v_{i}$ represents the value of vector $v$ at time $i$. We use bold notation to indicate sequences, i.e., $\mathbf{v}_{i}=[v_{1},\dots,v_{i}]$. We use $\sigma(\cdot)$ to denote the sigmoid function. The products $\odot$ and $\otimes$ denote the Hadamard (element-wise) product and the Kronecker (block-wise) product, respectively. $\mathbb{I}_{n}$ denotes the identity matrix of size $\mathbb{R}^{n\times n}$. Generally, we omit stating the bias term for weight matrices unless stating the bias term helps with clarity.
@@ 2 Preliminaries
In this section, we introduce the key architectural components studied in this work: attention, SSMs, and RNNs. We remark that these components are often the central block - considered to be the backbone - within a complex architecture composed of other blocks and skip connections (see for instance [13]). In what follows, we review exclusively the backbone block, which we denote as $f(\cdot)$ in $\mathbf{y}=f(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{L\times d}$ and $\mathbf{y}\in\mathbb{R}^{L\times d}$ are the input and output sequences, respectively.
@@ 2.1 Attention
The standard self-attention block [2] consists of three matrices: $W_{Q},\ W_{K},\ \text{and}\ W_{V}$, which are the learnt parameters of the model. These matrices, when multiplied with the input $\mathbf{u}$, yield the queries $\mathbf{q}\in\mathbb{R}^{d_{k}}$, keys $\mathbf{k}\in\mathbb{R}^{d_{k}}$, and values $\mathbf{v}\in\mathbb{R}^{d_{v}}$, respectively:
@@
where $\zeta(\cdot)$ is a map $\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}$ and is applied row-wise. For standard self-attention, the softmax function is used, i.e. $\zeta(\cdot)=\mathrm{softmax}(\cdot)$, but given the limitations of the softmax function, alternative formulations have been proposed. We consider two formulations of attention: softmax attention (2) and linear attention [14]. We focus on masked attention formulations, i.e., the attention matrix $\zeta(\mathbf{q}\mathbf{k}^{\top})$ has a lower-triangular structure, and to simplify the derivations, we drop the scaling factor $\sqrt{d_{k}}$.
@@ 2.2 State Space Models
Architectures based on a state space parametrization compute the output $\mathbf{y}$ through a dynamic recurrence of input signals at each time step $i$, i.e.,
@@
where $h_{i}$ is the hidden state of the system, and the dynamic matrices of appropriate dimensions $A_{i},B_{i},C_{i},D_{i}$ are the learnt model parameters. Different time-varying and time-invariant parameterizations for $A_{i},B_{i},C_{i},D_{i}$ have been proposed in the literature (an overview is given in [15]). Here we discuss the most prominent one.
@@
The first selective SSM parametrization (S6) was introduced together with the Mamba architecture [7]. The S6 block parametrizes the recurrence as
@@
with $\Delta_{i}=\mathrm{softplus}(W_{\Delta}(W_{u}u_{i})+b_{\Delta})$ for every $i$, $W_{\Delta}$, $W_{u}$, $W_{B}$, $W_{C}$, $W_{D}$, and $A$ are learnt matrices of appropriate dimensions, and $b_{\Delta}$ is a learnt bias. While SSM models allow for complex-valued matrices $A_{i},B_{i},C_{i},D_{i}$, here we restrict ourselves to real-valued matrices as in [7].
@@ 2.3 Recurrent Neural Networks
Similar to SSMs, RNNs also parameterize the input-output relationship via a recurrent computation, commonly given by the long short-term memory (LSTM) [16], i.e., at each time step $i$
@@
where $W_{f},W_{i},W_{o}$ and $U_{f},U_{i},U_{o}$ are the learnt gate parameters. In this paper, we focus on two variants: quasi LSTMs (qLSTM) [9], which removes the output dependence of the gates, and RG-LRU [10], which attempts to integrate ideas from SSMs into RNNs.
@@
The RG-LRU model presents a hybrid between a qLSTM and a SSM using the recurrence
@@ 3.1 Dynamical Systems Framework (DSF)
The DSF relies on a dynamical systems representation of the architectures. A dynamical system models how a system’s state, here denoted by $h$, evolves over time according to a difference or differential equation. Dynamical systems often evolve under the evolution of some input $u$, and the observable is an output $y$. These systems capture time-dependent processes, rendering them suitable for understanding the behavior of sequence models. Here, we choose a recurrent state space representation. This choice is motivated by the widespread use of state space model representations for dynamical systems. Moreover, we show in later sections that this representation encompasses attention, RNNs, and SSMs in a suitable fashion that allows for further analysis. In particular, a linear structured time-varying (LTV) dynamical system is defined by the recurrence
@@
where $h_{i}\in\mathbb{R}^{N}$ is the hidden state initialized with $h_{-1}=0$, $\Lambda_{i}\in\mathbb{R}^{N\times N}$ is the diagonal state transition matrix, $B_{i}\in\mathbb{R}^{N\times d}$ and $C_{i}\in\mathbb{R}^{d\times N}$ are the input and output matrices, respectively, and $D_{i}\in\mathbb{R}^{d\times d}$ is a scaled skip connection. Dynamical system (11) can alternatively be written in its convolutional representation, i.e., $\mathbf{y}=\mathbf{\Phi}\mathbf{u}$, where the convolutional kernel $\mathbf{\Phi}$ is defined as
@@
Note that the convolution kernel $\mathbf{\Phi}$ is of the same dimension as the attention matrix $\zeta(\mathbf{q}\mathbf{k}^{\top})$ and that these matrices are equivalent, up to the scaling factor $W_{V}$ used in self-attention.
@@
This recurrent view yields a causal convolution kernel by definition. However, certain models (e.g. non-masked attention) also use non-causal kernels. This can be incorporated in the DSF (11) by modifying the state update (11a). For the sake of simplicity and consistency with the recent literature, we stick with causal models in the following.
@@ 3.2 Architecture Reformulation
In the following, we show how popular architectures based on attention, SSMs, and RNNs can be rewritten into the DSF. To do this, all models will be reformulated into recurrence (11), i.e., all resulting DSF representations will have hidden state dimension $N$ [^1] The dimensions used in the following are visualized in Appendix A for clarity. Although the parametrization of models commonly found in the literature is conductive to efficient computation, here we depart from this convention. The goal of the DSF reformulation is to establish a theoretical framework that leads us to mathematical insights on the design of these models. The presented formulations are not intended to be computationally implemented in DSF form, however the framework can be used to identify computational algorithms for new architectures. For instance, the convolutional form of linear attention (12) is efficiently implemented via flash linear attention [17]. However, using the recurrent form derived below it can also be implemented via scan algorithms [18], e.g., parallel scan [5, 6] or accelerated scan [19]. Given that the structural requirements on the model parameterization of the algorithm is met, the DSF allows to identify existing algorithms to apply to new models even if the algorithm was designed for another model class.
@@ 3.2.1 Attention
In the following, we assume that we can separate the nonlinear map in (2) as
@@
where $\phi(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, $\psi(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, and $\eta(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m\times(i+1)}\rightarrow\mathbb{R}$, which is the case for all the considered architectures in this paper. Note that if $\zeta(\cdot)$ is a kernel function, the proposed separability is satisfied by construction, as it holds that $\phi=\psi$ and $\eta=1$. This allows us to write the self-attention input-output relationship as
@@
We note that for the recurrence (11), the matrix $\Lambda_{i}$ is given as an $nd\times nd$ matrix, where $n$ is the number of features in $\phi$ and $\psi$, and $d$ is the input dimension. However, due to the scalar structure of $\Lambda_{i}$ in (15a), it can be implemented as the scalar multiplication $\frac{\eta(q_{i-1},\mathbf{k}_{i-1})}{\eta(q_{i},\mathbf{k}_{i})}h_{i-1}$ in (11). Hence, the hidden state is never materialized as such in the computation of the attention scores. Interested readers are referred to Appendix B for a detailed derivation.
@@
In the case of linear attention, both maps $\phi(\cdot)$ and $\psi(\cdot)$ in the DSF parametrization (15) are separable and we use the kernel proposed in [14], i.e.,
@@
where $\textrm{elu}(\cdot)$ is the exponential linear unit.
@@
We also study generalized linear attention, where we require that the maps $\phi(\cdot)$, $\psi(\cdot)$ are linear, but allow for general nonlinear normalization functions $\eta(q_{i},\mathbf{k}_{i})$, i.e.,
$\phi(q_{i})=q_{i},\quad\psi(k_{j})=k_{j},\quad\eta(q_{i},\mathbf{k}_{i}).$ (17)
@@
Softmax attention also satisfies the assumption of separability (13). However, it holds that the feature vector representation of the transformed Gaussian kernel in the softmax function, i.e., $e^{q_{i}^{\top}k_{j}}$, is infinite dimensional. Hence, the DSF representation (15) of softmax attention (2) and its corresponding hidden state dimension $N$ would also be infinite dimensional. This insight gives further motivation to approximations of the softmax function by using, e.g., a Taylor series approximation such as in [20], to render the feature vector finite-dimensional.
@@
The exponential in softmax attention $e^{q_{i}^{\top}k_{j}}$ can be expressed in terms of its Taylor expansion, which consists of an infinite sum of polynomial kernels of increasing degree $p$, decomposable through the vectors of monomials $\bigotimes_{j=1}^{p}q_{i}$. See Appendix C for a complete proof. ∎
The work in [21] analyzes softmax attention as a kernel smoother and [14] shows that a kernel-based formulation can lead to linear complexity in sequence length for finite dimensional kernels. In [22], a kernel-based formulation of softmax is used to propose orthogonal random features to model softmax attention with linear complexity. In [20] a Taylor approximation of softmax attention is proposed, also leading to linear complexity. Finally, [23] relates transformer decoders to dynamical systems with increasing state size arising from the masked upper triangular part of the attention matrix. Compared to these works, we analyze how the proposed formulations compare in the recurrence (11) allowing us to compare to SSMs and RNNs in the following sections.